OpenStudy (anonymous):

Which of the following is slope intercept form of an equation? a. 2x + y = 7 b. 3y = - x + 7 c. y = 6x - 4 d. 3x - 2y = 7

OpenStudy (anonymous):

C is in Slope intercept form...

OpenStudy (anonymous):

thank you :) can u help me with more

OpenStudy (anonymous):

I can try lol

OpenStudy (anonymous):

Put the following equation into slope intercept form. 3x - 5y = 25 a. y = -3/5x + 5 b. y = 3/5 x - 5 c. y = -3/5x + 25 d. y = 3/5x + 25

OpenStudy (anonymous):

Just a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm not completely sure but i think that its either C or D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Which equation does not have a y-intercept of 5? a. 2x = y - 5 b. 3x + y = 5 c. y = x + 5 d. 2x - y = 5

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Which is an equation for the line with slope 1/3 through (-2, 1)? a. y = 1/3x + 1 b. y = 1/3x + 5/3 c. y = 1/3x - 5/3 d. y = 1/3x + 1/3

OpenStudy (kidrah69):

\[y-y1=m(x-x1)\]

OpenStudy (kidrah69):

\[y-1=1/3(x-(-2))\]

OpenStudy (kidrah69):

Thats the point slop formula needed to find the equation needed can you solve that?

OpenStudy (anonymous):

no

OpenStudy (kidrah69):

Ok whats \[1/3\times~X\]

OpenStudy (kidrah69):

remeber we are using the distributive property